OpenStudy (anonymous):

12+u/-5=-3 (u over -5)

OpenStudy (anonymous):

if you go to that website I can show you how to do it on a white board

OpenStudy (anonymous):

i cant leave this website not allowed , moms orders

OpenStudy (anonymous):

oh ok well u=3

OpenStudy (anonymous):

Just kidding that is wrong let me fix it

OpenStudy (anonymous):

Just kidding it is right u=3

OpenStudy (anonymous):

ARe you up for three more??

OpenStudy (anonymous):

ha ha sure

OpenStudy (anonymous):

8h+7=-113 33=5w-12 x/12-15=31

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Ok the first one you want to subtract 7 from both sides 8h+7=-113 8h=-120 divide each side by 8 h=-15

OpenStudy (anonymous):

Second problem you want to add 12 to both sides 33=5w-12 45=5w divide each side by 5 w=9

OpenStudy (anonymous):

third problem x/12-15=31 Is x over (12-15) or is x just over 12?

OpenStudy (anonymous):

x is just over 12

OpenStudy (anonymous):

ok so for the third problem x/12-15=31 multiply each side by 12 x-15=372 ad 15 to each side x=387

OpenStudy (anonymous):

I think those are right....

OpenStudy (anonymous):

The initial question, if it is 12 + [u/(-5)] = -3, then the answer is not u = 3

OpenStudy (anonymous):

You know what I think sklee is right lol sorry dumb mistake on my part U=75

OpenStudy (anonymous):

and the last problem x=552